produced by an injection-molding operation are checked for conformance to specifications. Each tool contains 14 cavities in which parts are produced, and these parts fell into a conveyor when the press opens. An inspector chooses 1 part(s) that do not conform to specifications. from among the 14 at random. Two cavities are affected by a temperature malfunction that results in parts (b) What is the probability that i the inspector finds at least one nonconforming part?
Answer #1

Here we have

sample size (n)= 1.

and probabilty of finding a defective is 2 malfuction cavity

so p=2/14

a) Now the  probabilty that inspector finds it defective =2/14 =0.1429

(b) Here sample size is 1 so the required probability will be same for both of the parts

i.e. p=2/14 =0.1429

.